### v2

# MetaphorShare: A Dynamic Collaborative Repository of Open Metaphor Datasets

Abstract
The metaphor studies community has developed numerous valuable labelled corpora in various languages over the years. Many of these resources are not only unknown to the NLP community, but are also often not easily shared among the researchers. Both in human sciences and in NLP, researchers could benefit from a centralised database of labelled resources, easily accessible and unified under an identical format. To facilitate this, we present MetaphorShare, a website to integrate metaphor datasets making them open and accessible. With this effort, our aim is to encourage researchers to share and upload more datasets in any language in order to facilitate metaphor studies and the development of future metaphor processing NLP systems. The website has four main functionalities: upload, download, search and label metaphor datasets. It is accessible at www.metaphorshare.com.
[…]
linecolor=black,linewidth=0.3pt
MetaphorShare: A Dynamic Collaborative Repository
of Open Metaphor Datasets
Joanne Boisson, Arif Mehmood  and Jose Camacho-Collados Cardiff NLP, School of Computer Science and Informatics Cardiff University, United Kingdom {boissonjc,mehmooda3,camachocolladosj}@cardiff.ac.uk
1 Introduction
The topic of figurative language processing has been addressed since early years of artificial intelligence research Martin (1990); Fass (1997); Kintsch (2000), inspired by philosophical Richards (1936), linguistic Wilks (1973) and then cognitive science theories Lakoff and Johnson (1980) and experimental studies Katz et al. (1988). In spite of the influence of the different disciplines on one another, the resources created to train and evaluate NLP models have often been distinct from the resources created for metaphor studies in other fields, with the notable exception of Master Metaphor List Lakoff (1994) and the VU Amsterdam Metaphor Corpus Steen (2010).
In recent years, metaphor processing has attracted more and more attention with the progress made possible by transformer-based language models, for example Mao et al. (2019) or Zeng and Bhat (2021), and in particular the large LMs such as GPT-3 Brown et al. (2020) used by Wachowiak and Gromann (2023) for metaphor identification. Several dedicated workshops with shared tasks Ghosh et al. (2022, 2024); Sharma et al. (2020); Tayyar Madabushi et al. (2022a) have been organized, leading to the creation of more resources. The proper handling of figurative language by models is of crucial importance for improving performance in downstream tasks Han et al. (2022); Li et al. (2024).
[…]
Figure 1: MetaphorShare search page. Specific datasets, languages, and tag types can be selected, and a text-based search within tagged expressions or into the entire text is implemented. Additional features provided with the record appear when clicking the Show Details button.
From the perspective of Humanities, tools for searching corpora such as Sketch Engine Kilgarriff et al. (2014), annotation tools Koller et al. (2008) and data analysis softwares have become part of the standard corpus-based research methodology. Effective adaptive support for metaphor identification and interpretation is the next step of this on-going inter-disciplinary collaboration. There were two limitations until recently: metaphor identification system were not accurate enough to be used on free text, and different researchers are looking for different metaphors.
On one side, numerous datasets labelled with metaphoric usage of words are created by researchers in many languages every year, and remain out of the scope of NLP research and evaluations of the systems because they are not publicly shared or consistently formatted. On the other, metaphor studies researchers often work on domain specific corpora, languages other than English (for which NLP has developed the most resources), and use different definitions or retain different types of metaphors for different projects, with no convenient way to systematically compare their labels with other researchers, or to rely on existing metaphor an automatic labelling tool that answer a project specific need. Moreover, these created datasets are not used by other communities or are often labeled using different schemata and formats.
We propose to help speeding up collaborations between AI/NLP communities and linguistics/metaphor studies by facilitating the unification of dataset formats and the access to existing resources by everyone, while preserving the information encoded in the original datasets. MetaphorShare is a website designed at sharing new and publicly available existing resources (see Figure 1). It is an online dynamic repository where anyone can upload or download open metaphor datasets. Our online labeling tool also offers the possibility to directly create a dataset, and optionally share it. Our hope is that the website will create a new synergy between different components of metaphor studies, limit the loss of datasets, catalog them and facilitate their automatic processing through a unified format. The targeted audience is anyone running corpus based research on metaphors, including NLP researchers for the development of metaphor processing and identification systems. [^1] An introduction to MetaphorShare is available at https://youtu.be/Fi48SOjueEE
2 Related work
Who studies and labels metaphors?
Cameron and Maslen (2010) summarizes the landscape of metaphor studies. Since Lakoff and Johnson (1980) Conceptual Metaphor Theory (CMT) core role in cognitive sciences, metaphors analysis have been done differently in linguistics and literature, but also became a standard methodology in Anthropology, Educational Research, Political Science or Management research. The purpose of metaphor analysis in these disciplines is to uncover latent meaning, sometimes elusive in corpora. Recently, Baleato Rodríguez et al. (2023) introduced propaganda modeling, extending the domain of use of NLP beyond isolated metaphor processing to the perspective social media content analysis.
Metaphor related websites.
Veale and Li (2012) develop a system of metaphor interpretation and generation, Metaphor Magnet, that relies on the harvest of stereotype from the Google n-grams. The stereotypical attributes associated to concepts are leveraged for the suggestion of relevant metaphors and their interpretation (e.g. given Google is –Microsoft, the system outputs giant with properties like lumbering and sprawling). More recently, Mao et al. (2023) release end-to-end metaphor identification and interpretation website for free text (non-domain specific): MetaPro tags metaphoric tokens or multi-word expressions in sentences, generates paraphrases, and outputs a more abstract concept mapping derived from the analysis of the sentence. The authors list limitations due to the lack of training data for inference and extended metaphors. The main difference with these two initiatives and ours is that our aim is to integrate already labeled datasets into a unified repository, encouraging the structuring of resources and expression of needs from the metaphor study community, before building automatic annotation tools. Similarly to ours, MetaPro also uses several metaphor datasets for training the models Mao et al. (2019), but in this case they do not mention more than six from the NLP community.
Annotation tools.
Many open source (e.g. Label Studio, Dataturks, Doccano Nakayama et al. (2018) and Potato Pei et al. (2022)) and proprietary (e.g. LabelBox, Prodigy or Amazon Mechanical Turk) annotation tools exist. Our purpose is not to compete with these services but rather to offer a simple targeted experience for metaphor annotation, in which the output is directly mapped to the unified repository format. This way, non-expert users can avoid the need for learning a general tool, as well as having to learn how to modify given labeled outputs, which wouldn’t be trivial for most.
Metaphor processing in NLP.
One important usage of our website is to facilitate the creation of personalised metaphor processing models. Metaphor processing in NLP comprises many methods developed for metaphor identification Turney et al. (2011); Tsvetkov et al. (2014); Mao et al. (2019); Wachowiak and Gromann (2023), but also generation Veale (2016); Stowe et al. (2021); Chakrabarty et al. (2021b), textual Mao et al. (2018) and multimodal Kulkarni et al. (2024) interpretation, metaphor understanding through entailment Agerri et al. (2008); Chakrabarty et al. (2021a); Stowe et al. (2022), among other tasks. Ge et al. (2023) provide a comprehensive recent survey on the topic. We do not currently store images in MetaphorShare, but the extension of the database to non textual data for multimodal research is straightforward.
Name Reference N %M
Psycholinguistics
JANK Jankowiak (2020) 24022footnotemark: 2 50
CARD_V Cardillo et al. (2010) 280 50
CARD_N Cardillo et al. (2010, 2017) 512 50
NLP
MOH Mohammad et al. (2016) 1632 25
NewsMet Joseph et al. (2023) 1205 49
TSV_A Tsvetkov et al. (2014) 1945 50
GUT Gutiérrez et al. (2016) 8591 54
Multi-Word Expressions
PVC Tu and Roth (2012) 1348 65
MAD Tayyar Madabushi et al. (2022b) 4558 48
MAGPIE Haagsma et al. (2020) 48395 75
MIPVU
VUAC_BO Boisson et al. (2023) 39223 52
TONG Tong et al. (2024) 1428 46
Table 1: Twelve metaphor datasets included in the MetaphorShare repository. The N column shows the number of instances in the datasets. %M indicates the percentage of expressions labelled as metaphorical.
3 Data Collection and Unification
MetaphorShare aims at facilitating computational research on multiple datasets. A minimally constrained format is required to store the datasets into a database, conveniently compare the resources, search efficiently into the different fields across datasets, and run experiments on several sets at once. Metaphor datasets can be added by their authors or by anyone if their license allows it. In this section, we review the existing dataset formats and our unification choices. The website database currently contains 25 datasets. We select twelve metaphor datasets with open licenses and diverse encoded information to illustrate our unification process and as representative examples for our repository. Their characteristics are summarized in Table 1 and in more details in Appendix Table 5.
One tagged expression per example.
Many datasets in NLP contain labels for binary classification or paraphrasing of one delimited expression within a given context. The expressions considered might be a single token, a multi-word expression (MWE), a phrase or a compound. The task of their original release paper may be to decide if the marked expression is used metaphorically or literally, or to find a correct literal paraphrase of the metaphoric expression, among other possible analysis and tasks. The context may consist in one single word, in the case of Adjective-Noun pairs datasets, as for example in Gutiérrez et al. (2016) or Tsvetkov et al. (2014). They often consist in one sentence such as Mohammad et al. (2016) or Cardillo et al. (2010, 2017) or Jankowiak (2020). Longer contexts are sometimes directly provided, such as in Tayyar Madabushi et al. (2022b) or Haagsma et al. (2020) who include two sentences before and after the sentence containing the marked Potential Idiomatic Expression (PIE), or indirectly provided with a reference to the original document, as for example in Joseph et al. (2023) who share references to newspaper articles or Tu and Roth (2012) who link the annotated examples to sentence pointers in the British National Corpus.
Multiple tags per example.
Steen (2010) defined an annotation guideline that is widely used in metaphor studies, the MIPVU procedure, together with the release of the VU Amsterdam Corpus (VUAC), the largest existing metaphor corpus. Other examples of English open datasets created following MIPVU, that are in MetaphorShare, are Nacey et al. (2019b) and Julich-Warpakowski (2022) on the Music Criticism domain. Many other MIPVU datasets exist in various languages with adaptations of the guideline Nacey et al. (2019a), for example Nacey (2022) for Scandinavian. In this framework, every token in a document is labelled as metaphoric or literal.
Different versions of the VUAC suiting NLP tasks, adapted for binary classification tasks Leong et al. (2018), or including additional ratings of novelty Parde and Nielsen (2018) have been released over the years. In the evaluation section of this paper, Section 5, we use a modified version of the VUAC corpus, adapted for binary classification (VUAC_BO) as described in Boisson et al. (2023).
Additional Information.
Beyond metaphorical/literal labels of expressions in context, metaphor datasets may contain more or less common additional information. There are scalar variables such as metaphoricity/figurativeness ratings Cardillo et al. (2010); Dunn (2014); Katz et al. (1988), confidence scores, concreteness scores, novelty scores Parde and Nielsen (2018), level of emotion scores Mohammad et al. (2016), frequency in corpora… They may also contain categorical variables such as PoS, textual source and genre, metaphor type, source and target concept/domains information Gordon et al. (2015). Figure 1 shows an example of the additional information encoded in the Mohammad et al. (2016) dataset for one instance, as diplayed on MetaphorShare. Some datasets, for example the LCC Mohler et al. (2016) and Gordon et al. (2015), also annotate the fragment of the context that carries the lexical information about the target of a metaphor (e.g. in the phrase ocean of happiness, the metaphoric expression is ocean and the target lexical cue of the metaphor is happiness).
Unified input format.
To simplify the search into various datasets and their handling for further system development, we require a specific input format for uploading a new resource on the website. MetaphorShare currently accepts CSV files having minimally one column named tagged_text. The column must contain text with at least one tagged expression per line. The accepted tag set is currently: < m: metaphoric, l: literal, t: target, a:semantically anomalous, u: free tag>. Each cell must minimally contain one tag and multiple tags are allowed, making possible the upload of MIPVU datasets without sentence duplication, e.g. I <m>swim</m> today in an <m>ocean</m> of <t>happiness</t>. Additional information may be added in freely named columns of the CSV file. Sentence indices can therefore be added to preserve the sentence order in a document and include metaphor in discourse datasets. Redundant fields across datasets such as PoS information have recommended names provided in the website guideline and we will see in Section 4.2 that the manual validation step by the website administrators allows a discussion with the user/author for further field names unification if needed. If any author of a dataset encounters issues to reformat a dataset into our expected format, we invite her or him to contact us by email for support.
Field Description
Required
Name Name of the person uploading data
Email Address of the person uploading data
Dataset Name of the dataset
Author Main author of the dataset
CSV file Formatted dataset
License New or existing dataset license
Optional
Paper title Publication presenting the dataset
Author(s) Authors of the paper
Year Paper’s publication year
Reference Bibtex
Languages Languages of the labelled examples
Source Source corpora of the labelled instances
Genre Novel, poetry, news, spoken language…
Target POS PoS of the target if any
Source POS PoS of the source expression if any
Annot. num. Number of annotators
Annot. profile Linguists, authors, crowdsourcing…
IAA Inter-annot. agreement metric & score
Comments Additional description and comments
Table 2: Required and optional dataset information to fill when uploading a new dataset
4 MetaphorShare: The Website
MetaphorShare general architecture is described in Section 4.1. The website is organized around four main pages: an uploading page (Section 4.2), a dataset catalog (Section 4.3), a page for searching into the datasets records (Section 4.4), and an online annotation tool (Section 4.5).
4.1 Website architecture
MetaphorShare is hosted by Cardiff University School of Computer Science and Informatics. It employs SSL protocols and HTTPS to safeguard data transmission. Direct access to the database is restricted, with backend schedulers managing all search and upload operations, thus fortifying data protection.
Back & front end.
The backend is developed with Python FastAPI, chosen for its speed and ease of use. It features schedulers for routine tasks such as data ingestion and file cleanup. Integration with Elasticsearch allows for efficient indexing and retrieval of dataset entries. Additionally, dataset metadata is securely stored in a PostgreSQL database, ensuring data integrity and structured storage. The frontend is developed using ReactJS (version 18), leveraging libraries such as Bootstrap and Ant Design. State management is handled by Redux, and data visualization is powered by Chart.js.
Database.
Our system employs a dual-database approach to efficiently manage and query data. A PostgreSQL database serves as the primary storage for user-submitted data during the dataset upload process, and stores them ensuring data integrity and facilitating efficient data management. An Elasticsearch engine enhances the search capabilities of MetaphorShare with exact and fuzzy match for text field supported in multiple languages.
Field Description
Required
id Database record ID
dataset_id Link to the Dataset index
expression Token or expression being labelled
label Metaphoric, literal, anomalous, other
position Expression offset in the given context
Optional (often present)
sent_index Index preserving the sentences order
reference Reference to a source document
pos Part of Speech
long_context Context surrounding a labelled sentence
target Lexical cue of the target concept
target_position target offset in the given context
source Metaphor source concept/domain
target Metaphor target concept/domain
mscore Metaphoricity/figurativeness score
Free Any field, for example emotion ratings
Table 3: Fields to describe an instance of a dataset in the ElasticSearch database
Train Test set (name & size)
J&C MOH NewsMet TSV_A GUT TONG PVC MAD MAGPIE VUAC_BO
132 732 305 1645 7691 528 448 47495 3658 38323
J&C .89 .52 .68 .78 .71 .54 .61 .63 .77 .60
MOH .70 .58 .42 .73 .67 .38 .21 .59 .26 .28
NewsMet .69 .45 .61 .72 .60 .38 .33 .48 .50 .38
TSV_A .66 .50 .61 .90 .83 .52 .39 .52 .61 .51
GUT .77 .40 .66 .73 .96 .61 .76 .65 .87 .68
TONG .39 .25 .14 .0 .0 .33 .53 .50 .03 .45
PVC .56 .41 .61 .57 .52 .54 .86 .62 .78 .57
MAD .71 .45 .54 .64 .64 .52 .79 .95 .82 .58
MAGPIE .71 .45 .67 .72 .74 .56 .73 .68 .93 .64
VUAC_BO .55 .41 .65 .73 .77 .59 .55 .61 .79 .67
Table 4: Results of the cross dataset evaluation. F1-score of the metaphor class.
4.2 Uploading a new dataset
Once a dataset is formatted into a CSV file, as described in Section 3, and after its license is specified by the authors, the dataset can be uploaded by any user without registering.
Dataset metadata.
The required and optional fields to be filled in at uploading time are presented in Table 2. For example, if a dataset is associated with one or several publications, the references should be added in the appropriate fields, preferably in the BibTeX format. The information provided will be displayed in the dataset catalog (c.f. Figure 2 in the Appendix), and will help to further categorize the datasets in the future versions of the website.
Automatic and manual validations.
After validating the form, an immediate automatic file format check on the data is done and feedback is returned to the user for acceptance or rejection of the file. Outputs of this step can be seen in Appendix Figure 3 & 4. In a second manual validation step, the website administrators check the license information, dataset reference, and eventually suggest modifications of CSV free columns names, that contain the dataset specific additional information, in an email interaction with the user, for database field unification purposes.
Storage in the database.
Once a dataset passes the two validation steps, it is stored in the main repository database. This ElasticSearch database has two indices: the Datasets index similar to Table 2 fields stores all the user provided metadata, and information extracted during the processing of the data file such as the dataset size and the label distribution. The second one, the Potentially Metaphoric Expressions, PME index, contains the output of the data file parse, as shown Table 3. Each textual labelled example is linked to the Dataset index.
4.3 Dataset Catalog
The main purpose of this page is to display the main features of each dataset included in the repository. The fields filled at uploading time, including the license chosen, are displayed in the dataset catalog page. Statistics computed from the input file are also shown and plotted, in particular the number of instances, label set and distribution. A partial view of a dataset summary in the catalog is shown in the Appendix, Figure 2.
4.4 Search Page
The records can be accessed and downloaded through a search page, as shown in Figure 1. The search fields include optional selections of datasets and languages from drop-down lists, restricted search for a label (e.g. metaphorical or literal), and keyword search into the PME text (with exact and fuzzy match), or into the entire text. One phrase or sentence is shown by default on the result page, but longer context can be accessed in one click, as well as all the additional information attached to a record. [^2] The organization of the catalog and the fine-grained search options are likely to evolve in future versions of MetaphorShare. The result of any search can be downloaded as a CSV file.
4.5 Online data annotation
On the Edit/Label Dataset page, one can choose to directly input some text into the interface of the annotation tool, or alternatively to upload a table (in a CSV or Excel format) containing unannotated or partially annotated text. An input can then be labelled from scratch or edit labels by pairing highlighted expressions with tags (c.f. Figure 5 in the Appendix). While the MetaphorShare tagset is predefined, custom tags can also be created. The resulting labelled data (c.f. Appendix Figure 6) can then be downloaded as a CSV, or directly uploaded in MetaphorShare following the procedure described in Section 4.2. This online annotation functionality facilitates the creation of resources immediately compatible with our accepted formats, and speeds up the correction of CSV files that do not pass the automatic validation step of the dataset upload procedure. It is also an accessible tool for annotators without any computer science background.
5 Evaluation
As a sanity check, we first verified that MetaphorShare uploading, database insertion and search functionalities are working properly with the datasets presented in Table 2.
Then, as a case study we perform a cross-dataset metaphor identification analysis to illustrate a possible usage of MetaphorShare for NLP research. Given an input expression, the task consists of deciding whether the expression is a metaphor or not. Different projects label metaphors for different purposes. MetaphorShare makes it easy to fine-tune models on relevant dataset to further support project-specific automatic or semi-automatic labeling.
Experimental setting.
In order to maintain a training set of the same size for all datasets, we randomly sampled 800 examples from each set to create the training sets. Because our datasets from psycholinguisics are small and similar in their creation method, we grouped them into one single set called J&C in this experiment. RoBERTa base models Liu et al. (2019) are then finetuned independently on the 10 datasets and tested on each of them. Similarly to the experiments in Boisson et al. (2023), for hyperparameter optimisation, we rely on the Bayesian Optimization with the Hyperband (BOHB) algorithm Falkner et al. (2018) available in RayTune Liaw et al. (2018), with an identical hyperparameter search space and 25 trials.
Results.
Results are shown in Table 4. As expected, models trained and evaluated on the same dataset often achieved the best results. A few datasets generalise better than others. It is surprisingly the case for J&C and GUT, both containing only short example with constrained syntactic structures. On the other hand, datasets such as TONG and NEWS do not generalize as well, probably due to different labelled PoS, to the token or MWE span of the labels, or due to redundant contexts corresponding to datasets created for other tasks such as paraphrasing in the case of TONG.
6 Conclusion and Future Work
In this paper, we have presented a website to label, unify and share metaphor datasets. It enables an easy integration of new resources of different original formats by the community. The 25 currently included datasets are in English and we are planning to integrate and reach communities working on other languages in the future as one of the main aims.
In the future we will focus on integrating a functionality for automatically labelling raw text. The progress recently made by transformer-based language models for metaphor identification opens the path to the creation of helpful personalized tools for automatic or semi-automatic labelling of metaphors based on an initial sample of annotations.
Acknowledgments
This research was supported by the Association for Researching and Applying Metaphors through the Building Bridges fund. We thank the British National Corpus for giving us the permission to share fragments of texts of the BNC directly on MetaphorShare. We also thank Eileen Cardillo, all the authors of Gutiérrez et al. (2016) and of Birke and Sarkar (2006) for allowing us to redistribute their datasets online for research purposes, and to the rest of the authors of the datasets included in the repository for the creation of metaphor datasets with a flexible open license. Without the effort of the metaphor community, this work would not have been possible.
Appendix A Detailed information of datasets upoaded in MetaphorShare and used in our experiments
Table 5 provides detailed information about the datasets, including license, source corpora, label distribution and sentence distribution. Additional links and preprocessing details are listed below :
JANK:
Anomalous sentences are not shown in the table.
MOH:
The license is available at https://saifmohammad.com/WebPages/metaphor.html. The original dataset contains 1639 instances. A few duplicated example sentences caused by orthographic variants of the target word, such as distil/distill have been removed.
NewsMet:
We show the manually labelled sentences (named gold by the authors). The corpus from which the sentences are extracted can be found at https://github.com/several27/FakeNewsCorpus.
TSVET_A:
GUT:
The UKaC is presented in Ferraresi et al. (2008).
PVC:
The original dataset is accessible at https://cogcomp.seas.upenn.edu/page/resource_view/26
MAGPIE:
The PMB corpus from which some sentences are sourced is presented in Abzianidze et al. (2017).
TONG:
Original VUAC sentences and apt (dataset new label) paraphrases are counted in the table.
Appendix B Dataset upload
Automatic file format check.
Figure 3 & 4 show the row by row feedback provided after the automatic validation step when the file is rejected, and the pop up message indicating that a file passed the automatic validation step.
Dataset information after validation.
Figure 2 shows a part of a dataset presentation page once it is integrated into the catalog and database.
Additional Information attached to a labelled example after validation.
Appendix C Dataset catalog
Figure 2 is a screenshot of the top of the page presenting a dataset in the Catalog section of MetaphorShare. The dataset name, main author and research paper associated to the data release are displayed.
Appendix D Online annotation tool
Figure 5 shows the annotation tool interface with one example sentence. Figure 6 shows how the created records of a dataset are displayed, allowing an easy step by sep construction of a dataset.
[…]
Figure 2: Screenshot of top of the the datasets information page in the catalog section of the website. The English dataset released with Jankowiak (2020) is presented as an example.
[…]
Figure 3: Screenshot of the file format check for a rejected file. The line the error occurs in the CSV file and the type of errors are specified.
[…]
Figure 4: Screenshot of the file format check after a CSV file is accepted for manual review.
[…]
Figure 5: Screenshot of the online annotation tool showing the text input area, tag selection and creation, and coloured resulting tagged text.
[…]
Figure 6: Screenshot showing dataset rows available for tagging or edition, as displayed in the annotation tool.